What is the interior angle of a regular 16 sided polygon by using the formula (n-2)*180? Can you write the answer step by step please.
Answer by Alan3354(69443)

What is the interior angle of a regular 16 sided polygon by using the formula (n-2)*180?
----------------
n = 16
The total of the 16 Int angles is (16-2)*180 = 2520 degs
Each angle = 2520/16 = 157.5 degs
---------------
Do these this way:
The sum of exterior angles for ALL polygons is 360 degs
360/16 = 22.5 degs Ext
180 - 22.5 = 157.5 degs Int
---------
Less work, no formula to remember.